OpenStudy (anonymous):

A pole 3 m high has a shadow 5 m long when the shadow of a nearby building is 110 m long. How tall is the building? The answer should be to the nearest tenth or as a whole number. x = ?

OpenStudy (cwrw238):

use property of similar triangles - corresponding sides are in same ratio.

OpenStudy (anonymous):

First lets call the pole (P), the pole's shadow PS, and the building's shadow BS, and finally we will call the building B Now we just need to set up a proportion: P/PS = B/BS 3/5 = B/110 cross multiplying, we get: 330 = 5B B = 66m tall.

OpenStudy (anonymous):

wow that made it easier for me to understand :) thanks a lot